Shadee Corp. expects to sell 560 sun visors in May and 360 in June. Each visor sells for $17. Shadees beginning and ending fhi i was hoping someone could help me solve this problem, ive been on it for hours and can see to get the answer right
Answer #1

Calculate total cash receipts

August September
cash Sales 8500*60% = 5100 7650*60% = 4590
Collection from credit sales
July Sales 9010*40%*35% = 1261
August sales 8500*40%*50% = 1700 1190
September sales 7650*40%*50% = 1530
Total Cash receipt 8061 7310